A child of mass 30 kg travels in a lift calculate the force exerted on the child by the floor of the lift when the lift moves 1.downward with an acceleration of 1 metre per second 2.upward with an acceleration of 1 metre per second what is the weight of the child when the lift falls freely. g =10ms^-2
1
Expert's answer
2021-10-06T11:45:47-0400

1.

"F=mg-ma=30(10-1)=270" N


2.

"F=mg+ma=30(10+1)=330" N


When the lift falls freely, a=g, so:

"F=mg-mg=0"

This means that the weight of the child is zero.